# 1D $Z_{2}$ lattice gauge theory in periodic Gauss law sectors

### v1

@@ Abstract
We calculate the properties of a 1D $Z_{2}$ lattice gauge theory in different Gauss law sectors, corresponding to different configurations of static charges, set by the orientations of the gauge spins. Importantly, in quantum simulator experiments these sectors can be accessed without adding any additional physical particles or changing the Hamiltonian: The Gauss law sectors are simply set by the initial conditions. We study the interplay between conservation laws and interactions when the static charges are chosen to form periodic patterns. We classify the different Gauss law sectors and use the density matrix renormalization group to calculate the ground state compressibility, density profiles, charge density wave order parameters, and single particle correlation functions as a function of matter density. We find confined and deconfined phases, charge density waves, correlated insulators, and supersolids.
@@ I Introduction
Non-uniform Gauss law sectors have received very little attention. The one notable exception is the concept of “disorder-free localization” where initial states prepared in a superposition of different Gauss law sectors experience an effective binary disorder potential. This randomness localizes the matter particles, despite the fact that the Hamiltonian is translationally invariant [11, 12, 13, 14, 15].
@@ II Model
${H}=-\sum_{i}(b_{i+1}^{{\dagger}}\sigma^{z}_{i,i+1}b_{i}+h.c.)-h\sum_{i}\sigma^{x}_{i,i+1}.$ (1)
Here $b,b^{\dagger}$ are annihilation and creation operators of hardcore bosons on the sites, $i$ is the site index and $\sigma^{x},\sigma^{z}$ are Pauli operators for the spin-1/2 particles on links. When a boson hops, it flips the spin on the link. This is the $Z_{2}$ analog of minimal coupling where one interprets $\sigma_{z}=e^{i\pi/2(\sigma_{z}-1)}=e^{iA_{1}}$, where $A_{1}$ is the spatial component of the vector potential. In this interpretation the electric field, which is conjugate to the vector potential, is $E=(\sigma_{x}+1)/2$. The electric field takes on the values $0,1$, and the Zeeman term can be interpreted as the energy associated with this electric field. In this formulation there is no time-like component of the vector potential, as one assumes the temporal gauge $A_{0}=0$.
@@ III Results
Writing the two vacuum states as $A$ and $B$, the spin state with two bosons will have the form $ABA$, with two domain walls. When the vacuum magnetization is finite, the energy density of the $B$ region is larger than the $A$ region, binding pairs together. This phenomenon has been well described in the uniform sector [7]. The energy of a pair is linear in the separation between the bosons. When the vacuum magentization vanishes, this pairing mechanism is absent.
This pairing is analogous to Quark confinement in QCD, and these paired phases are referred to as confined. A simple example is contrasting the +++++++++++ + + + + and ------ - - - - sectors. In the absence of particles the former has a net magnetization, while the latter does not. Thus for infinitesmally low densities one expects them to respectively be confined and deconfined. These sectors are, however, connected by a particle-hole transformation. Thus as the density approaches $1$ we expect a deconfined phase in the +++++++++++ + + + + sector.
@@
Figure 3: The charge gap, $\Delta=(E(N+2)+E(N-2)-2E_{N})/2$ extrapolated to $1/L\to 0$ at various rational boson fillings, $n_{b}$ for the Gauss law patterns, $G^{(1)}=+-+-+-$ (red dashed) and $G^{(2)}=+--+--$ (blue solid) at $h=1.5$ and $h=1$ respectively. The largest gaps are off-scale: $\Delta\approx 4h,6h$ at $n_{b}=1/2,2/3$ for $G^{(1)}$ and $G^{(2)}$.
Before giving our results for the single particle correlations we first argue for the presence of incompressible states at particular densities. We numerically search for these insulating states by calculating the ground state energy of $N$ particles in a system of length $L$, $E_{L}(N)$. The chemical potential in the thermodynamic limit is $\mu=E_{L}(N+1)-E_{L}(N)$. An incompressible (insulating) state is characterized by a discontinuity in the chemical potential, characterized by the gap $\Delta=\lim_{L\to\infty}(E_{L}(N+2)+E_{L}(N-2)-2E_{N})/2$. We use the finite Density Matrix Renormalization Group (DMRG) to calculate $E_{L}(N)$ for several different values of $L$, and then extrapolate to the thermodynamic limit by taking the intercept as $1/L\to 0$. Figure 3 shows representative examples, which will be discussed in detail below. As is known from previous work [7], in the uniform sector +++++++++++ + + + + or its particle-hole transform, ------ - - - -, the only insulators are the trivial ones where there is zero or one particle per site.
In Fig. 3 we see gapped states at various rational values of $n_{b}$. The largest gaps are seen when the density of particles is equal to $n_{m}$, the fraction of sites with $G_{i}=-1$. We refer to the ratio $\nu=n_{b}/n_{m}$ as the filling. In Sec. III.1 we discuss the integer filling insulators (commensurate insulators), where $\nu=1$. In Sec. III.2 we discuss the fractional filling insulators (incommensurate insulators), where $\nu$ is a rational number with a non-trivial denominator.
@@ III.1.1 Instability towards a commensurate charge density wave
Figure 6: Magnitude of the charge density wave order parameter, $O_{cdw}=2/L\sum_{i}(-1)^{i}n^{b}_{i}$ as a function of field $h$ in the Gauss law pattern, $G=+-+-+-$ using DMRG for a system size $L=100$ at half filling. As $h\to 0$, the order parameter’s slope diverges.
@@ III.3 Single particle correlations
Figure 10: Single particle off-diagonal correlations, $C_{ij}=\langle a^{\dagger}_{i}a_{j}\rangle$ are fit to an exponential function ($C_{ij}=e^{-|i-j|/\xi}$). The plot shows $\xi$ as a function of boson density, $n_{b}$ for the uniform Gauss law sector, $G=++++$ for bond dimension, $\chi=20,30,40,50$ using the VUMPS infinite DMRG algorithm. Correlation length $\xi$ appears to diverge only as $n_{b}\to 1$
To characterize the compressible phases, we calculate the single-particle density matrix $\langle a_{i}^{\dagger}a_{j}\rangle$. Generically we expect the bosons to be confined, in which case the correlations fall off exponentially at large distance, $\langle a^{{\dagger}}_{i}a_{j}\rangle=a_{0}e^{-|i-j|/\xi}$. To calculate this correlation function, we make an infinite matrix product state ansatz for the ground state wavefunction, and use the VUMPS algorithm [21] in the Itensor library to optimize the parameters. By directly working in the thermodynamic limit we avoid any boundary effects, but the bond dimension $\chi$ of the ansatz introduces an effective cutoff [23]. Thus we repeat the calculation with multiple $\chi$, and study how $\xi$ evolves. For this calculation, we use the Gauss’s law to eliminate the bosons and get a local Hamiltonian of just the gauge spins on the links. We don’t explicitly impose boson number conservation and instead use a chemical potential term to obtain an average boson density in the ground state (see Appendix B for calculation details). Thus for each value of $\mu$ we calculate both $\xi$ and $n_{b}$.
Figure 10 shows the behavior of the single particle correlation length in the uniform sector, $G=+++++$. As can be seen, for any density $n_{b}\neq 1$, the correlation length is finite, indicating that the particles are confined. The correlation function diverges as $n_{b}\to 1$, indicating that the holes are deconfined for infinitesmal hole density. In the graph we show the results from several different bond dimensions, illustrating convergence for all but the highest densities.
@@
We find similar results in other sectors: For generic $n_{b}\neq 0,1$, we always find a finite correlation length. The correlation length will diverge as $n_{b}\to 0$ if we are in a net zero magnetiztion sector. Similarly, it diverges as $n_{b}\to 1$ if the net magnetization vanishes under a particle-hole transformation. For example in the alternating section $G=+-+-+-$, the correlation length diverges both as $n_{b}\to 0$ and $n_{b}\to 1$. We never see a diverging correlation length at the charge-density wave transitions. This latter observation is sensible, as the effective model for excess particles/holes on top of the charge density waves always has a net magnetization.
@@ III.4 Supersolidity
For generic filling we have a compressible state with no off-diagonal long range order in the single-particle density matrix. Analogous to what has been seen in previous works, the two-particle density matrix, however, has long-range (power law) order, and the system is a pair superfluid [7]. Additionally it is energetically favorable for particles to sit on sites where $G_{i}=-1$ to minimzs the Zeeman energy of the gauge field spins. Thus $O_{cdw}$ defined in Sec. III.1 is non-zero. Simultaneously having pair superfluidity and charge-density-wave order could be described as supersolidity. One caveat is that the periodicity is imposed by the conserved quantities. Nonetheless the Hamiltonian is translationally invariant.
@@ IV Summary and Outlook
Experiments have begun to explore simple lattice gauge theories. One important feature is that by changing their initial conditions they can access a range of different ‘Gauss law sectors’, which have patterns of stationary charges. Prior work has almost exclusively focused on the uniform sector. We study ground state phases in periodic Gauss law sectors. We classify them by the net magnetization of the gauge fields in the ground state in the absence of matter particles. A small (infinitesmal) number of particles added to a finite magnetization sector will bind up into pairs, while in a zero magnetization sector they remain unpaired. These can be interpreted as confined and deconfined phases. One can apply a particle-hole transformation and use equivalent reasoning to study a small (infninitesmal) number of holes – finding both confined and deconfined states. By numerically studying the single particle density matrix we find that as long as the density is not equal to 00 or $1$ the particle/holes always pair and are confined, regardless of sector.
@@ Appendix B Spin model for calculating correlation functions
We will formulate this transformation as a mapping between operators that act on the spin-particle space ($\sigma^{x}_{i,i+1},n_{i}$) and a spin chain ($\tau^{x}_{i}$). For notational simplicity, in the pure spin model, we place the spins on sites, rather than bonds. In the $x$-basis, the spin wavefunction is found by simply ignoring the particles, and hence $\tau_{i}^{x}=\sigma_{i,i+1}^{x}$. Flipping a single spin changes the boson parity on the adjacent sites, and hence $\tau_{i}^{z}=(b_{i}+b_{i}^{\dagger})\sigma_{i,i+1}^{z}(b_{i+1}+b_{i+1}^{\dagger})$. Flipping a string of spins corresponds to adding or removing a single particle $\prod_{j\geq i}\tau_{j}^{z}=(b_{i}+b_{i}^{\dagger})\prod_{j\geq i}\sigma^{z}_{j,j+1}=a_{i}+a_{i}^{\dagger}$, where the operators $a_{i}$ and $a_{i}^{\dagger}$ were introduced in Sec. II when we eliminated the spins. They correspond to the simplest gauge invariant ways of adding or removing particles. The operator $a_{i}$ vanishes on a state where no particles are on site $i$. Thus $a_{i}=(a_{i}+a_{i}^{\dagger})(1-(-1)^{n_{i}})/2$. We can then use Gauss’ law, $G_{i}=\sigma^{x}_{i-1,i}(-1)^{n_{i}}\sigma^{x}_{i,i+1}$, to conclude
@@
$\begin{split}\hat{H}-\mu\hat{N}=\frac{1}{2}\sum_{i}\left(\tau^{x}_{i-1}G_{i}G_{i+1}\tau^{x}_{i+1}-1\right)\tau^{z}_{i}\\ -h\sum_{i}\tau_{i}^{x}+\frac{\mu}{2}\sum_{i}\tau^{x}_{i-1}G_{i}\tau^{x}_{i}\end{split}$ (18)

### v2

# One-dimensional $Z_{2}$ lattice gauge theory in periodic Gauss-law sectors

@@ Abstract
We calculate the properties of a one-dimensional $Z_{2}$ lattice gauge theory in different Gauss law sectors, corresponding to different configurations of static charges, set by the orientations of the gauge spins. Importantly, in quantum simulator experiments these sectors can be accessed without adding any additional physical particles or changing the Hamiltonian: The Gauss law sectors are simply set by the initial conditions. We study the interplay between conservation laws and interactions when the static charges are chosen to form periodic patterns. We classify the different Gauss law sectors and use the density matrix renormalization group to calculate the ground state compressibility, density profiles, charge density wave order parameters, and single particle correlation functions as a function of matter density. We find confined and deconfined phases, charge density waves, correlated insulators, and supersolids.
@@ I Introduction
Non-uniform Gauss law sectors have received very little attention. The one notable exception is the concept of “disorder-free localization” where initial states prepared in a superposition of different Gauss law sectors experience an effective binary disorder potential. This randomness localizes the matter particles, despite the fact that the Hamiltonian is translationally invariant [11, 12, 13, 14, 15, 16].
@@ II Model
${H}=-t\sum_{i}(b_{i+1}^{{\dagger}}\sigma^{z}_{i,i+1}b_{i}+h.c.)-h\sum_{i}\sigma^{x}_{i,i+1}.$ (1)
Here $b,b^{\dagger}$ are annihilation and creation operators of hardcore bosons on the sites, $i$ is the site index and $\sigma^{x},\sigma^{z}$ are Pauli operators for the spin-1/2 particles on links. The lattice spacing $a$ is set to unity. Throughout this work, we set $t=1$. When a boson hops, it flips the spin on the link. This is the $Z_{2}$ analog of minimal coupling where one interprets $\sigma_{z}=e^{i\pi/2(\sigma_{z}-1)}=e^{iA_{1}}$, where $A_{1}$ is the spatial component of the vector potential. In this interpretation the electric field, which is conjugate to the vector potential, is $E=(\sigma_{x}+1)/2$. The electric field takes on the values $0,1$, and the Zeeman term can be interpreted as the energy associated with this electric field. In this formulation there is no time-like component of the vector potential, as one assumes the temporal gauge $A_{0}=0$.
@@ III Results
Writing the two vacuum states as $A$ and $B$, the spin state with two bosons will have the form $ABA$, with two domain walls. When the vacuum magnetization is finite, the energy density of the $B$ region is larger than the $A$ region, binding pairs together. This phenomenon has been well described in the uniform sector [7]. The energy of a pair is linear in the separation between the bosons. When the vacuum magnetization vanishes, this pairing mechanism is absent.
This pairing is analogous to Quark confinement in QCD, and these paired phases are referred to as confined. A simple example is contrasting the +++++++++++ + + + + and ------ - - - - sectors. In the absence of particles the former has a net magnetization, while the latter does not. Thus for infinitesimally low densities one expects them to respectively be confined and deconfined. These sectors are, however, connected by a particle-hole transformation. Thus as the density approaches $1$ we expect a deconfined phase in the +++++++++++ + + + + sector.
@@
Figure 3: The dimensionless charge gap, $\Delta=(E(N+2)+E(N-2)-2E_{N})/2t$ (with $t=1$) extrapolated to $1/L\to 0$ at various rational boson fillings, $n_{b}$ for the Gauss law patterns, $G^{(1)}=+-+-+-$ (red dashed) and $G^{(2)}=+--+--$ (blue solid) at $h=1.5$ and $h=1$ respectively. The largest gaps are off-scale: $\Delta\approx 4h,6h$ at $n_{b}=1/2,2/3$ for $G^{(1)}$ and $G^{(2)}$.
Before giving our results for the single particle correlations we first argue for the presence of incompressible states at particular densities. We numerically search for these insulating states by calculating the ground state energy of $N$ particles in a system with $L$ sites, $E_{L}(N)$. The chemical potential in the thermodynamic limit is $\mu=E_{L}(N+1)-E_{L}(N)$. An incompressible (insulating) state is characterized by a discontinuity in the chemical potential, characterized by the dimensionless gap $\Delta=\lim_{L\to\infty}(E_{L}(N+2)+E_{L}(N-2)-2E_{N})/2t$. We use the finite Density Matrix Renormalization Group (DMRG) to calculate $E_{L}(N)$ for several different values of $L$, and then extrapolate to the thermodynamic limit by taking the intercept as $1/L\to 0$. Figure 3 shows representative examples, which will be discussed in detail below. As is known from previous work [7], in the uniform sector +++++++++++ + + + + or its particle-hole transform, ------ - - - -, the only insulators are the trivial ones where there is zero or one particle per site.
In Fig. 3 we see gapped states at various rational values of $n_{b}$ where $n_{b}=N/L$. The largest gaps are seen when the density of particles is equal to $n_{m}$, the fraction of sites with $G_{i}=-1$. We refer to the ratio $\nu=n_{b}/n_{m}$ as the filling. In Sec. III.1 we discuss the integer filling insulators (commensurate insulators), where $\nu=1$. In Sec. III.2 we discuss the fractional filling insulators (incommensurate insulators), where $\nu$ is a rational number with a non-trivial denominator.
@@ III.1.1 Instability towards a commensurate charge density wave
Figure 6: Magnitude of the charge density wave order parameter, $O_{cdw}=2/L\sum_{i}(-1)^{i}n^{b}_{i}$ as a function of field $h/t$ in the Gauss law pattern, $G=+-+-+-$ using DMRG for a system size $L=100$ sites at half filling. As $h/t\to 0$, the order parameter’s slope diverges.
@@ III.3 Single particle correlations
Figure 10: Single particle off-diagonal correlations, $C_{ij}=\langle a^{\dagger}_{i}a_{j}\rangle$ are fit to an exponential function ($C_{ij}=e^{-|i-j|/\xi}$). The plot shows dimensionless correlation length (measured as number of sites) $\xi$ as a function of boson density, $n_{b}$ for the uniform Gauss law sector, $G=++++$ for bond dimension, $\chi=20,30,40,50$ using the VUMPS infinite DMRG algorithm. Correlation length $\xi$ appears to diverge only as $n_{b}\to 1$
To characterize the compressible phases, we calculate the single-particle density matrix $\langle a_{i}^{\dagger}a_{j}\rangle$. Generically we expect the bosons to be confined, in which case the correlations fall off exponentially at large distance, $\langle a^{{\dagger}}_{i}a_{j}\rangle=a_{0}e^{-|i-j|/\xi}$. Here $\xi$ is a dimensionless correlation length measured as number of sites. To calculate this correlation function, we make an infinite matrix product state ansatz for the ground state wavefunction, and use the VUMPS algorithm [22] in the Itensor library to optimize the parameters. By directly working in the thermodynamic limit we avoid any boundary effects, but the bond dimension $\chi$ of the ansatz introduces an effective cutoff [24]. Thus we repeat the calculation with multiple $\chi$, and study how $\xi$ evolves. For this calculation, we use the Gauss’s law to eliminate the bosons and get a local Hamiltonian of just the gauge spins on the links. We don’t explicitly impose boson number conservation and instead use a chemical potential term to obtain an average boson density in the ground state (see Appendix B for calculation details). Thus for each value of $\mu$ we calculate both $\xi$ and $n_{b}$.
Figure 10 shows the behavior of the single particle correlation length in the uniform sector, $G=+++++$. As can be seen, for any density $n_{b}\neq 1$, the correlation length is finite, indicating that the particles are confined. The correlation function diverges as $n_{b}\to 1$, indicating that the holes are deconfined for infinitesimal hole density. In the graph we show the results from several different bond dimensions, illustrating convergence for all but the highest densities.
@@
We find similar results in other sectors: For generic $n_{b}\neq 0,1$, we always find a finite correlation length. The correlation length will diverge as $n_{b}\to 0$ if we are in a net zero magnetization sector. Similarly, it diverges as $n_{b}\to 1$ if the net magnetization vanishes under a particle-hole transformation. For example in the alternating section $G=+-+-+-$, the correlation length diverges both as $n_{b}\to 0$ and $n_{b}\to 1$. We never see a diverging correlation length at the charge-density wave transitions. This latter observation is sensible, as the effective model for excess particles/holes on top of the charge density waves always has a net magnetization.
@@ III.4 Supersolidity
For generic filling we have a compressible state with no off-diagonal long range order in the single-particle density matrix. Analogous to what has been seen in previous works, the two-particle density matrix, however, has long-range (power law) order, and the system is a pair superfluid [7]. Additionally it is energetically favorable for particles to sit on sites where $G_{i}=-1$ to minimize the Zeeman energy of the gauge field spins. Thus $O_{cdw}$ defined in Sec. III.1 is non-zero. Simultaneously having pair superfluidity and charge-density-wave order could be described as supersolidity. One caveat is that the periodicity is imposed by the conserved quantities. Nonetheless the Hamiltonian is translationally invariant.
@@ IV Summary and Outlook
Experiments have begun to explore simple lattice gauge theories. One important feature is that by changing their initial conditions they can access a range of different ‘Gauss law sectors’, which have patterns of stationary charges. Prior work has almost exclusively focused on the uniform sector. We study ground state phases in periodic Gauss law sectors. We classify them by the net magnetization of the gauge fields in the ground state in the absence of matter particles. A small (infinitesimal) number of particles added to a finite magnetization sector will bind up into pairs, while in a zero magnetization sector they remain unpaired. These can be interpreted as confined and deconfined phases. One can apply a particle-hole transformation and use equivalent reasoning to study a small (infinitesimal) number of holes – finding both confined and deconfined states. By numerically studying the single particle density matrix we find that as long as the density is not equal to 00 or $1$ the particle/holes always pair and are confined, regardless of sector.
@@ Appendix B Spin model for calculating correlation functions
We will formulate this transformation as a mapping between operators that act on the spin-particle space ($\sigma^{x}_{i,i+1},n_{i}$) and a spin chain ($\tau^{x}_{i}$). For notational simplicity, in the pure spin model, we place the spins on sites, rather than bonds. In the $x$-basis, the spin wavefunction is found by simply ignoring the particles, and hence $\tau_{i}^{x}=\sigma_{i,i+1}^{x}$. Flipping a single spin changes the boson parity on the adjacent sites, and hence $\tau_{i}^{z}=(b_{i}+b_{i}^{\dagger})\sigma_{i,i+1}^{z}(b_{i+1}+b_{i+1}^{\dagger})$. Note that this term includes both a nearest-neighbor hopping and a pair creation/annihilation of bosons. Flipping a string of spins corresponds to adding or removing a single particle $\prod_{j\geq i}\tau_{j}^{z}=(b_{i}+b_{i}^{\dagger})\prod_{j\geq i}\sigma^{z}_{j,j+1}=a_{i}+a_{i}^{\dagger}$, where the operators $a_{i}$ and $a_{i}^{\dagger}$ were introduced in Sec. II when we eliminated the spins. They correspond to the simplest gauge invariant ways of adding or removing particles. The hard core boson’s annihilation operator $a_{i}$ vanishes on a state where no particles are on site $i$. Thus $a_{i}=(a_{i}+a_{i}^{\dagger})(1-(-1)^{n_{i}})/2$. We can then use Gauss’ law, $G_{i}=\sigma^{x}_{i-1,i}(-1)^{n_{i}}\sigma^{x}_{i,i+1}$, to conclude
@@
$\begin{split}\hat{H}-\mu\hat{N}=-\frac{t}{2}\sum_{i}\left(\tau^{x}_{i-1}G_{i}G_{i+1}\tau^{x}_{i+1}-1\right)\tau^{z}_{i}\\ -h\sum_{i}\tau_{i}^{x}+\frac{\mu}{2}\sum_{i}\tau^{x}_{i-1}G_{i}\tau^{x}_{i}\end{split}$ (18)